Given the linear equation:
- Find the y-coordinates of the given points. Show your work.
- Plot those points and graph the linear equation.

y = 1/3 x - 2

(0, ), (3, ), (-3, )

Found 2 solutions by texttutoring, rapaljer:
Answer by texttutoring(324)

Just substitute the value into x, and solve for y.
i) x=0
y=1/3(0) -2
y=0-2
y=-2

(0,-2) is a point.

ii) x=3
y=(1/3)*3 - 2
y=1-2
y=-1

(3,-1) is another point.
iii) y=(1/3)*(-3) - 2
y=-1-2
y=-3

(-3,-3) is the last point.

Plot those three points on a graph, and then connect the dots to draw the line. If you want to see a picture of this, let me know.

Answer by rapaljer(4671)

y=1/3 x - 2

If x=0,
y =0-2
y =-2

If x=3,
y =1/3*3-2
y =1-2
y =-1

If x=-3,
y=1/3*(-3) -2
y=-1-2
y=-3

Dr. Robert J. Rapalje, Retired